OpenStudy (anonymous):

The resistivity of gold is 2.44×10-8 Ω•m at room temperature. A gold wire that is 1.9 mm in diameter and 33 cm long carries a current of 250 mA. What is the electric field in the wire?

OpenStudy (fifciol):

\[E=\frac{ V }{ l}=\frac{ IR }{ l }\] \[R=\frac{ \rho l }{ A }\] \[E=\frac{ I \rho l }{Al }=\frac{ 4I \rho }{\pi d^2 }\]

OpenStudy (anonymous):

what does p, stand for

OpenStudy (fifciol):

it's rho, resistivity

OpenStudy (anonymous):

rho as,p=charge / area,

OpenStudy (fifciol):

p=RA/l

OpenStudy (anonymous):

so if I want to substitute from your equation, instead of p, what should i put

OpenStudy (fifciol):

resistace * area / length

OpenStudy (anonymous):

okay let me check, ill get back to you

OpenStudy (anonymous):

so the equation will end up to be E=IR/L

OpenStudy (fifciol):

yes, but you have given resistivity rho, not resistance R

OpenStudy (anonymous):

do I have to change the unit of the current to be A not mA, if it so how will I do that

OpenStudy (fifciol):

1 A = 10^3 mA

OpenStudy (anonymous):

look this are the possible answers that I have .2×10−3 V/m 5.4×10−3 V/m 2.7×10−3 V/m 5.4×10−4 V/m 1.7×10−3 V/m

OpenStudy (anonymous):

sorry buddy, I have to go there is something that I have to do. but please can you do it for me and when you are done post it for me

OpenStudy (fifciol):

I find 2.15 * 10-3